Write down an equation to represent the production A, X, B, C

1
Expert's answer
2022-03-07T18:08:01-0500

The standard form of a linear equation is written as Ax+By=CA+By

The standard form is useful in calculating the linear full equtions

linear function is an algebraic equation in which each term is either a constant or the product of a constant and (the first power of) a single variable. For example, a common equation, y

=mx+b

y=mx+b

(namely the slope-intercept form, which we will learn more about later) is a linear function because it meets both criteria with x

x

 and yy

 as variables and m

m and bb

 as constants. It is linear: the exponent of the xx

 term is a one (first power), and it follows the definition of a function: for each input (xx) there is exactly one output (yy). Also, its graph is a straightline.